OpenStudy (anonymous):

solve over domain -3pi/2 greater or equal to x < pi/2 for (cos2x)(cosx) - (sin2x)(sinx) = 0

OpenStudy (anonymous):

divide by cosx althrought what do we have

OpenStudy (anonymous):

\[\cos2x-\sin2xtanx=0\]

OpenStudy (anonymous):

cos 3x = 0

OpenStudy (anonymous):

\[\cos2x-2sinxcosx.sinx/cosx\]

OpenStudy (anonymous):

\[\cos2x-2\sin ^{2}x=0\]

OpenStudy (anonymous):

\[\cos ^{2}x-\sin ^{2}x-2\sin ^{2}x=0\]

OpenStudy (anonymous):

cos (alpha + beta) trig identities

OpenStudy (anonymous):

not so comfortable with that refer you to @inyourhead

OpenStudy (anonymous):

I don't do math buddy.

OpenStudy (anonymous):

cool @inyourhead that was the only name i can spell out so i havent met you on any Qs what is up